# Dual Language Models: Balancing Training Efficiency and Overfitting Resilience

### v1

@@ Abstract
This paper combines autoregressive and masked-diffusion training objectives without any architectural modifications, resulting in flexible language models that outperform single-objective models. Autoregressive modeling has been a popular approach, partly because of its training efficiency; however, that comes at the cost of sensitivity to overfitting. On the other hand, masked-diffusion models are less efficient to train while being more resilient to overfitting. In this work, we demonstrate that dual-objective training achieves the best of both worlds. To derive the optimal ratio between both objectives, we train and evaluate 50 language models under varying levels of data repetition. We show that it is optimal to combine both objectives under all evaluated settings and that the optimal ratio is similar whether targeting autoregressive or masked-diffusion downstream performance.
@@ 1 Introduction
The dominant paradigm for training large language models has been autoregressive next-token prediction (NEURIPS2020_1457c0d6). This approach is remarkably efficient in training, allowing models to quickly absorb vast amounts of text. However, this efficiency comes with a significant drawback: a tendency to overfit, especially when training data is limited or repeated (NEURIPS2023_9d89448b). This issue is becoming increasingly critical as the community reaches the so-called data wall – the imminent exhaustion of high-quality text data required to train ever-larger models according to established scaling laws (10.5555/3692070.3694094).
An alternative approach, masked-diffusion language modeling, offers a compelling solution to the overfitting problem. These models are inherently more robust to data repetition and can learn powerful bidirectional representations (prabhudesai2025diffusionbeatsautoregressivedataconstrained; ni2025difflm). Yet, this robustness comes at the cost of lower training efficiency; masked-diffusion models are known to be 16 times less sample-efficient than their autoregressive counterparts (nie2025scaling). The opposing failure modes of these two approaches – autoregressive models overfit while masked-diffusion models underfit – suggest a natural solution: their combination could yield the benefits of both.
@@
In this work, we show that it is possible to achieve the best of both worlds by simultaneously training a single language model on both autoregressive and masked-diffusion objectives. The core idea is to use the training efficiency of the autoregressive objective for rapid initial learning while using the masked-diffusion objective to regularize the model and prevent it from overfitting. The effectiveness of this dual-objective approach is illustrated in Figure 1. In the extreme data-constrained setting with 128 data repetitions, the purely autoregressive model learns quickly but then catastrophically overfits. The masked-diffusion model is immune to overfitting but converges very slowly. Our proposed dual-objective model combines the strengths of both and successfully leverages the given compute and data.
Building on this observation, we conduct a large-scale systematic study to find the optimal balance between these two objectives under varying degrees of data constraint. Our primary contributions are:
@@
  • Through an extensive empirical study, we systematically map the relationship between data repetition, the ratio of training objectives, and final downstream performance.
  • We demonstrate that a dual-objective approach is superior to single-objective training in all evaluated settings, for both autoregressive and masked-diffusion evaluation – including the finding that dual models outperform pure masked-diffusion models even in regular data settings.
  • We derive two practical recommendations for setting the optimal objective ratio when training in both regular and data-constrained regimes, providing a concrete guideline for future training of large language models.
  • We show that the dual language models can generalize to prefix language models at inference time, which further increases their downstream performance.
@@ 2 Background
As the name suggests, language models are statistical models $p_{\bm{\theta}}(\cdot)$ of the true language distribution of some training corpus $\mathcal{D}$. The training corpus consists of sequences $\bm{x}=(x_{1},x_{2}\dots x_{N})\in\mathcal{D}$ of subword tokens. The language models are trained by finding such parameters $\bm{\theta}$ that maximize the likelihood estimation (MLE; fisher1922mathematical; fisher1925theory):
$\mathop{\text{argmax}}_{\bm{\theta}}\mathop{\mathbb{E}}_{\bm{x}\,\sim\,\mathcal{D}}\Bigl[\log p_{\bm{\theta}}(\bm{x})\Bigr].$ (1)
In this paper, we combine two popular approaches for computing $p_{\bm{\theta}}(\cdot)$, autoregressive language models and masked-diffusion language models.
@@ 2.1 Autoregressive language modeling
Language models have a long tradition and since their inception in the seminal paper by shannon1951prediction, they have been factored into a chain of next-token prediction terms $p_{\bm{\theta}}(x_{i}\,|\,\bm{x}_{<i})$:
$\log p_{\bm{\theta}}(\bm{x})=\sum_{i=1}^{|\bm{x}|}\log p_{\bm{\theta}}(x_{i}\mid\bm{x}_{<i}).$ (2)
Computation of the next-token likelihoods can be efficiently parallelized when modeled by transformer networks (NIPS2017_3f5ee243), and thanks to their scalability, it has been the most popular paradigm behind the recent era of large language models (NEURIPS2020_1457c0d6).
@@ 2.2 Masked-diffusion language modeling
Masked-diffusion language models have recently become a popular alternative to autoregressive models (NEURIPS2021_958c5305; 10.5555/3692070.3693403; 10.5555/3737916.3742051; ou2025your; nie2025large). Computing $p_{\bm{\theta}}(\cdot)$ with masked-diffusion is slightly more complicated than with autoregression, but the resulting language model learns to handle full bidirectional context, which can lead to increased performance on downstream tasks (berglund2024the; 10.5555/3737916.3738000).
First, following NEURIPS2021_958c5305, we define the forward (and backward) diffusion process that gradually turns a sequence of tokens $\bm{x}$ into special mask tokens (and vice-versa). The diffusion process $\left\{\bm{x}^{t}\right\}$ depends on the time variable $t\in[0,1]$ so that $\bm{x}^{(0)}=\bm{x}$ and $\bm{x}^{(1)}$ is a fully masked sequence. The intermediate values are defined by the probability distribution $q$:
@@
We can see that each token can either remain unchanged or turn into a mask token with probability $t$. The forward process is fully reversible and we can accordingly define the backward process, which gradually unmasks a sequence (NEURIPS2021_958c5305). Using the results from ou2025your, the probability distribution $q_{0|t}(x_{i}|\,\bm{x}^{t})$ governing the backward process can be modeled with a time-independent transformer language model with parameters $\bm{\theta}$ as $p_{\bm{\theta}}(x_{i}\,|\,\bm{x}^{t})$. This model can be fitted to the training data by maximizing the lower bound on the log-likelihood estimate (ou2025your):
$\log p_{\bm{\theta}}(\bm{x})\geq\int_{0}^{1}\mathop{\mathbb{E}}_{\bm{x}^{t}\sim q_{t\mid 0}(\cdot\mid\bm{x})}\Biggl[\frac{1}{t}\sum_{\left\{i\mid x_{i}^{t}=\,\texttt{mask}\right\}}\!\!\!\!\!\!\!\!\!\log p_{\bm{\theta}}(x_{i}\mid\bm{x}^{t})\Biggr]\,\mathrm{d}t.$ (4)
The integral can be equivalently written as the expectation over $t\sim\mathcal{U}(0,1)$, thus, it can be directly used as a training objective when estimated by Monte-Carlo sampling (metropolis1949monte). Such a Monte-Carlo estimate can also be used at inference-time for likelihood-based evaluation, similarly to Equation 2. Note that the resulting objective is very similar to the one used to train masked language models such as BERT (devlin-etal-2019-bert).
3 Dual language modeling
The method of combining autoregressive and masked (diffusion) objectives is mostly based on the earlier GPT-BERT approach by charpentier-samuel-2024-bert. They showcased promising results for very small language models trained within the limitations of the BabyLM Challenge (conll-2024-babylm). We extend their approach to masked-diffusion language models and to orders of magnitude larger computation scale.
Dual objective and next-token prediction
Our goal is to align the two factorizations of the MLE objective in Equations 2 and 4 so that they can be parameterized by a single transformer model. For this reason, we use a slightly modified version of masked language modeling called masked next-token prediction (MNTP; lv-etal-2024-analysis). With this approach, the model always uses the hidden state at position $i$ to predict the next token at position $i+1$ (we prove that this parameterization is as expressive in Appendix E). In this way, both modes of operation are unified as they both perform next-token prediction, as illustrated in Figure 2. MNTP has also been used in recent work for adapting a masked diffusion model from an autoregressive checkpoint (gong2025scaling; ye2025dream7bdiffusionlarge).
@@ 3 Dual language modeling
Loss weighting
It is crucial to correctly weight the masked-diffusion objective by $\nicefrac{{1}}{{t}}$, as in Equation 4, to maintain the lower bound. Thus, on average, the masked diffusion objective is weighted by $\mathop{\mathbb{E}}_{t\sim\mathcal{U}(0,1)}[\nicefrac{{1}}{{t}}]=2$. To address this imbalance regarding the autoregressive objective, we double the weight of the autoregressive loss.
GPU-wise objective separation
In practice, naively mixing both objectives within a single batch could result in reduced throughput. For this reason, we assign each GPU device to a single objective so that the computation graph remains simple and static, and can be efficiently compiled. To be specific, we distribute the training of each model across 256 devices, which allows for choosing between 256 ratios of diffusion and autoregressive training. For example, if we wanted each global batch to contain as many diffusion samples as autoregressive samples, we would refer to this setting as the $1:1$ AR-D ratio (autoregressive-diffusion). A standard autoregressive model would be trained with $1:0$ AR-D ratio, and a model heavily skewed towards the masked-diffusion objective would be trained with $1:255$ AR-D, for example.
@@ 4 Evaluation
While it is common practice to only consider the value of loss on a held-out set when evaluating language models (kaplan2020scalinglawsneurallanguage; 10.5555/3600270.3602446; NEURIPS2023_9d89448b), it is important to measure the actual downstream performance to accurately assess the effect of different training configurations. This is especially crucial when training with two incompatible training losses.
@@
ARC-Easy (ARC-E) 2 3762\,376 $4^{\dagger}$ test clark2018thinksolvedquestionanswering
ARC-Challenge (ARC-C) 1 1721\,172 $4^{\dagger}$ test clark2018thinksolvedquestionanswering
BLiMP 67 00067\,000 22 warstadt-etal-2020-blimp-benchmark
Commonsense QA (CSQA) 1 2211\,221 55 val talmor-etal-2019-commonsenseqa
HellaSwag (HSwag) 10 04210\,042 44 val zellers-etal-2019-hellaswag
MMLU 14 04214\,042 44 test hendrycks2021measuring
OpenBook QA (OBQA) 500500 44 test mihaylov-etal-2018-suit
Physical Interaction QA (PIQA) 1 8381\,838 22 val Bisk_2020
Social IQa (SIQA) 1 9541\,954 33 val sap-etal-2019-social
@@
We follow the guidelines of the OLMES paper (gu-etal-2025-olmes) for the normalization of our log-likelihood estimations as well as the prompt format, with two changes: 1) we only evaluate in a zero-shot fashion to simplify the setup, 2) we only consider their “cloze” formulation of each task, which is more suitable for smaller models. For the BLiMP task, which is not considered in the OLMES evaluation suite, we do not apply any length normalization and take the raw log-likelihood score. Since the BLiMP and MMLU tasks contain multiple sub-tasks (67 for BLiMP, and 57 for MMLU), we report their macro-average as the final score. More information on how each task is normalized can be found in Appendix B.
@@
To ensure a fair aggregation of the different task scores, we first normalize the scores such that the random baseline of each task is at 0 and the maximum is at $1$; similarly to the Open LLM Leaderboard (open-llm-leaderboard-v2). To achieve this we apply the following formula to our scores: $\operatorname{score}(x,t)=\nicefrac{{(x-r_{t})}}{{(m_{t}-r_{t})}}$, where $x$ is the raw score, $r_{t}$ is the random baseline and $m_{t}$ is the optimal score for task $t$. We then take the simple average of the normalized scores across all tasks as the final performance of our model.
@@ 4.2 Masked-diffusion (bidirectional) evaluation
One possibility to evaluate the masked-diffusion capabilities of our models is to also leverage the training objective in Equation 4 and estimate the conditional log-likelihood of each completion by Monte-Carlo sampling. We describe this approach in more detail in Appendix C. While it provides accurate downstream scores, it is computationally expensive and less accurate than using simpler pseudo log-likelihood (PLL; wang-cho-2019-bert; salazar-etal-2020-masked; 10.5555/3737916.3738000) estimation.
PLL allows us to do bidirectional evaluation more than ten times faster while being more accurate than Monte-Carlo sampling (Appendix H). Therefore, we use PLL for evaluating the bidirectional capability of our models. We fully describe this method in Appendix D. As visualized in Figure 3 on the left, we specifically use the semi-autoregressive variation of PLL proposed by 10.5555/3737916.3738000.

### v2

# Dual-objective Language Models: Training Efficiency Without Overfitting

@@ Abstract
This paper combines autoregressive and masked-diffusion training objectives without any architectural modifications, resulting in flexible language models that outperform single-objective models. Autoregressive modeling has been a popular approach, partly because of its training efficiency; however, that comes at the cost of sensitivity to overfitting. On the other hand, masked-diffusion models are less efficient to train while being more resilient to overfitting. In this work, we demonstrate that dual-objective training achieves the best of both worlds. To derive the optimal balance between both objectives, we train and evaluate 50 language models under varying levels of data repetition. We show that it is optimal to combine both objectives under all evaluated settings and that the optimal balance is similar whether targeting autoregressive or masked-diffusion downstream performance.
@@ 1 Introduction
The dominant paradigm for training recent language models has been autoregressive next-token prediction (Brown et al., 2020). This approach is remarkably efficient in training, allowing models to quickly absorb vast amounts of text. However, this comes with a significant drawback: a tendency to overfit when training data is repeated (Muennighoff et al., 2023). This issue is becoming increasingly critical as the community reaches the so-called data wall – the imminent exhaustion of available data required to train ever-larger models according to established scaling laws (Villalobos et al., 2024).
An alternative approach, masked-diffusion language modeling, offers a compelling solution to the overfitting problem (Prabhudesai et al., 2025; Ni et al., 2025). Yet, this robustness comes at the cost of lower training efficiency; masked-diffusion models are known to be much less sample-efficient than their autoregressive counterparts (Nie et al., 2025a). The opposing failure modes of these two approaches – autoregressive models overfit while masked-diffusion models underfit – suggest a natural solution: their combination could yield the benefits of both.
@@
In this work, we show that it is possible to achieve the best of both worlds by simultaneously training a single language model on both autoregressive and masked-diffusion objectives. The core idea is to use the training efficiency of the autoregressive objective for rapid initial learning while using the masked-diffusion objective to regularize the model and prevent it from overfitting. The effectiveness of this dual-objective approach is illustrated in Figure 1. In the extreme data-constrained setting with 128 data repetitions, the purely autoregressive model learns quickly but then catastrophically overfits. The masked-diffusion model is immune to overfitting but converges very slowly. Our proposed dual-objective model combines the strengths of both and successfully leverages the given compute and data. The resulting models can be deployed as a standard autoregressive models with no inference overhead.
Building on this observation, we conduct a large-scale systematic study to find the optimal balance between these two objectives under varying degrees of data constraints. Our main contributions are:
@@
  • Through an extensive empirical study, we systematically map the relationship between data repetition, the ratio of training objectives, and final downstream performance. Demonstrating that our dual-objective approach is superior to single-objective training in all evaluated settings, for both autoregressive and masked-diffusion evaluation – including the finding that dual-objective models outperform pure masked-diffusion models even in regular data settings.
  • We derive two practical recommendations for setting the optimal objective ratio when training in both regular and data-constrained regimes, providing concrete guidelines for future training of large language models.
@@ 2 Background
Before diving into details of combining autoregressive and masked-diffusion models, we need to briefly describe those two modeling approaches and language modeling in general. As the name suggests, language models are statistical models $p_{\bm{\theta}}(\cdot)$ of the true language distribution of some training corpus $\mathcal{D}$. The training corpus consists of sequences $\bm{x}=(x_{1},x_{2},\dots x_{N})\in\mathcal{D}$ of subword tokens. The language models are trained by finding such parameters $\bm{\theta}$ that maximize the likelihood estimation (MLE; Fisher, 1922; 1925):
$\mathop{\text{argmax}}_{\bm{\theta}}\mathop{\mathbb{E}}_{\raisebox{-0.3pt}{${}_{\bm{x}\,\sim\,\mathcal{D}}$}}\Bigl[\log p_{\bm{\theta}}(\bm{x})\Bigr].$ (1)
In this paper, we combine two popular approaches for computing $p_{\bm{\theta}}(\cdot)$, autoregressive language modeling and masked-diffusion language modeling.
@@ 2.1 Autoregressive language modeling
Language models have a long tradition and since their inception in the seminal paper by Shannon (1951), they have been factored into a chain of next-token prediction terms $p_{\bm{\theta}}(x_{i}\,|\,\bm{x}_{<i})$:
[…]
Computation of the next-token likelihoods can be efficiently parallelized when modeled by transformer networks (Vaswani et al., 2017), and thanks to their scalability, it has been the most popular paradigm behind the recent era of large language models (Brown et al., 2020).
@@ 2.2 Masked-diffusion language modeling
Masked-diffusion language models have recently become a popular alternative to autoregressive models (Austin et al., 2021; Lou et al., 2024; Sahoo et al., 2025; Ou et al., 2025; Nie et al., 2025b). Computing $p_{\bm{\theta}}(\cdot)$ with masked-diffusion is slightly more complicated than with autoregression, but the resulting language model learns to handle full bidirectional context, which can lead to increased performance on downstream tasks (Berglund et al., 2024; Samuel, 2025).
First, following Austin et al. (2021), we define the forward (and backward) diffusion process that gradually turns a sequence of tokens $\bm{x}$ into special mask tokens (and vice-versa). The diffusion process $\left\{\bm{x}^{t}\right\}$ depends on the time variable $t\in[0,1]$ so that $\bm{x}^{(0)}=\bm{x}$ and $\bm{x}^{(1)}$ is a fully masked sequence. The intermediate values are defined by the probability distribution $q$:
@@
We can see that each token can either remain unchanged or turn into a mask token with probability $t$. The forward process is fully reversible and we can accordingly define the backward process, which gradually unmasks a sequence (Austin et al., 2021). Using the results from Ou et al. (2025), the probability distribution $q_{0|t}(x_{i}|\,\bm{x}^{t})$ governing the backward process can be modeled with a time-independent transformer language model with parameters $\bm{\theta}$ as $p_{\bm{\theta}}(x_{i}\,|\,\bm{x}^{t})$. This model can be fitted to the training data by minimizing the upper bound on the negative log-likelihood estimate (Ou et al., 2025):
[…]
The integral can be equivalently written as the expectation over $t\sim\mathcal{U}(0,1)$, thus, it can be directly used as a training objective when estimated by Monte-Carlo sampling (Metropolis and Ulam, 1949). Such a Monte-Carlo estimate can also be used at inference-time for likelihood-based evaluation, similarly to Equation 2. Note that the resulting objective is very similar to the one used to train masked language models such as BERT (Devlin et al., 2019).
3 Dual-objective language modeling
The method of combining autoregressive and masked (diffusion) objectives is mostly based on the earlier GPT-BERT approach by Charpentier and Samuel (2024). They showed promising results for very small language models trained within the limitations of the BabyLM Challenge (Hu et al., 2024). We extend their approach to masked-diffusion language models and to orders of magnitude larger computation scale.
Dual-objective language models are trained by minimizing the following combined loss function, which is further explained below in more detail:
$\mathop{\text{argmin}}_{\bm{\theta}}\mathop{\mathbb{E}}_{\raisebox{-0.3pt}{${}_{\bm{x}\,\sim\,\mathcal{D}}$}}\Bigl[{\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{AR}}(}\bm{x}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9};}\,\bm{\theta}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9})}+{\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}(1-\alpha)}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{MD}}(}\bm{x}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9};}\,\bm{\theta}{\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9})}\Bigl].$ (5)
@@ 3 Dual-objective language modeling
Loss weighting
The balance between the autoregressive objective $\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{AR}}$} and masked-diffusion objective $\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{MD}}$} is controlled by the hyperparameter $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$}. It is crucial for controlling the trade-off between training efficiency and overfitting robustness; its relation to the number of data repetitions is extensively tested by the following experiments.
[…]
Diffusion as next-token prediction
Our goal is to align $\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{AR}}$} and $\color[rgb]{0.1,0.1,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.1,0.1,0.9}\mathcal{L}_{\text{MD}}$} so that they can be parameterized by a single transformer model. For this reason, we use a slightly modified version of masked language modeling called masked next-token prediction (MNTP; Lv et al., 2024). With this approach, the model always uses the hidden state at position $i$ to predict the next token at position $i+1$ (we prove that this parameterization is as expressive as the standard approach in Appendix F). In this way, both modes of operation are unified as they both perform next-token prediction, as illustrated in Figure 2. MNTP has also been used in recent work for adapting a masked diffusion model from an autoregressive checkpoint (Gong et al., 2025; Ye et al., 2025).
@@ 4 Evaluation
While it is common practice to only consider the value of loss on a held-out set when evaluating language models (Kaplan et al., 2020; Hoffmann et al., 2022; Muennighoff et al., 2023), it is important to measure the actual downstream performance to accurately assess the effect of different training configurations. This is especially crucial when training with two incompatible training losses. That being said, we also report validation losses in Appendix G.
@@
ARC-Easy (ARC-E) 2 3762\,376 $4^{\dagger}$ test Clark et al. (2018)
ARC-Challenge (ARC-C) 1 1721\,172 $4^{\dagger}$ test Clark et al. (2018)
BLiMP 67 00067\,000 22 Warstadt et al. (2020)
Commonsense QA (CSQA) 1 2211\,221 55 val Talmor et al. (2019)
HellaSwag (HSwag) 10 04210\,042 44 val Zellers et al. (2019)
MMLU 14 04214\,042 44 test Hendrycks et al. (2021)
OpenBook QA (OBQA) $500$ 44 test Mihaylov et al. (2018)
Physical Interaction QA (PIQA) 1 8381\,838 22 val Bisk et al. (2020)
Social IQa (SIQA) 1 9541\,954 33 val Sap et al. (2019)
@@
We follow the guidelines of the OLMES paper (Gu et al., 2025) for the normalization of our log-likelihood estimations as well as the prompt format, with two changes: 1) we only evaluate in a zero-shot fashion to simplify the setup, 2) we only consider their “cloze” formulation of each task, which is more suitable for smaller models. For the BLiMP task, which is not considered in the OLMES evaluation suite, we do not apply any length normalization and score each sample with the raw log-likelihood score. Since the BLiMP and MMLU tasks contain multiple sub-tasks (67 for BLiMP, and 57 for MMLU), we report their macro-average as the final score. More information on how each task is normalized can be found in Appendix C.
@@
To ensure a fair aggregation of the different task scores, we first normalize the scores such that the random baseline of each task is at 0 and the maximum is at $1$; similarly to the Open LLM Leaderboard (Fourrier et al., 2024). To achieve this we apply the following formula to our scores: $\operatorname{score}(x,t)=\nicefrac{{(x-r_{t})}}{{(m_{t}-r_{t})}}$, where $x$ is the raw score, $r_{t}$ is the random baseline and $m_{t}$ is the optimal score for task $t$. We then take the simple average of the normalized scores across all tasks as the final performance of our model.
@@ 4.2 Masked-diffusion (bidirectional) evaluation
One possibility to evaluate the masked-diffusion capabilities of our models is to also leverage the training objective in Equation 4 and estimate the conditional log-likelihood of each completion by Monte-Carlo sampling. We describe this approach in more detail in Appendix D. While it provides accurate downstream scores, it is computationally expensive and less accurate than using simpler pseudo log-likelihood (PLL; Wang and Cho, 2019; Salazar et al., 2020; Samuel, 2025) estimation.
PLL allows us to do bidirectional evaluation more than ten times faster while being more accurate than Monte-Carlo sampling (Appendix J). Therefore, we use PLL for evaluating the bidirectional capability of our models. We fully describe this method in Appendix E. As visualized in Figure 3 on the left, we specifically use the semi-autoregressive variation of PLL proposed by Samuel (2025).